what is the perimeter of a regular hexagon whose sides are 52cm
Found 2 solutions by Nate, venugopalramana:
Answer by Nate(3500)

hexagon is a 6-sided figure ....
52 + 52 + 52 + 52 + 52 + 52 = 6 * 52 = 312cm

Answer by venugopalramana(3286)

what is the perimeter of a regular hexagon whose sides are 52cm
REGULAR HEXAGON HAS 6 EQUAL SIDES
SINCE EACH SIDE IS 52 CM,PERIMETER =6*52=312 CM